### v1

@@ Abstract
Ophthalmologists often rely on multimodal data to improve diagnostic precision. However, data on complete modalities are rare in real applications due to a lack of medical equipment and data privacy concerns. Traditional deep learning approaches usually solve these problems by learning representations in latent space. However, we highlight two critical limitations of these current approaches: (i) Task-irrelevant redundant information existing in complex modalities (e.g., massive slices) leads to a significant amount of redundancy in latent space representations. (ii) Overlapping multimodal representations makes it challenging to extract features that are unique to each modality. To address these, we introduce the Essence-Point and Disentangle Representation Learning (EDRL) strategy that integrates a self-distillation mechanism into an end-to-end framework to enhance feature selection and disentanglement for robust multimodal learning. Specifically, Essence-Point Representation Learning module selects discriminative features that enhance disease grading performance. Moreover, the Disentangled Representation Learning module separates multimodal data into modality-common and modality-unique representations, reducing feature entanglement and enhancing both robustness and interpretability in ophthalmic disease diagnosis. Experiments on ophthalmology multimodal datasets demonstrate that the proposed EDRL strategy outperforms the state-of-the-art methods significantly. Code is available at GitHub Repository.
@@
footnotetext: *These authors contribute equallyfootnotetext: Corresponding authors: Zongyuan.Ge@monash.edu, Imran.razzak@mbzuai.ac.ae
@@ 1 Introduction
To this end, we propose Essence-point and Disentangle Representation Learning (EDRL) framework. Specifically, the Essence-point Representation Learning (EPRL) module models the essence-points that guide the selection of discriminative information within each modality, effectively reducing task-irrelevant redundant representation. For modality feature disentanglement, the Disentangle Representation Learning (DiLR) module decomposes feature embeddings into modality-common and modality-unique components. It aligns shared information by driving the cross-correlation matrix toward a unit form while ensuring that modality-specific features remain distinct by approaching a zero form. Furthermore, we incorporate a self-distillation mechanism with two pipelines, one with missing modality inputs and another with complete modalities. Using the complete pipeline as a reference, this mechanism enhances missing information reconstruction, improving robustness under incomplete-modality conditions. Consequently, the EDRL framework minimizes redundancy, reduces inter-modality overlap, and enhances discriminative, interpretable, and generalizable multimodal learning. We evaluated the proposed method on intra and inter-modality incompleteness conditions across three ophthalmology multimodal datasets, where our approach achieves state-of-the-art performances.
Overall, our contributions are threefold:
  • We propose EPRL framework to reduce intra-modal redundancy by modeling prototypes for discriminative instance selection and incorporating self-distillation mechanism.
  • We introduce DiLR to disentangle multimodal features into independent modality-unique and modality-common representation.
  • Experiments on three ophthalmology multimodal datasets demonstrate the effectiveness of the EDRL strategy.
@@ 2.1 Problem Formulation
In this work, we propose an EDRL framework to address the challenge of missing modalities, particularly for solving (1) Inter-modality missing, where one modality is completely missing, and (2) Intra-modality missing, where natural noises are added into the modalities. To reduce "Task-irrelevant Redundant Information", we propose Essence-Point Representation Learning (EPRL) module to select the task-relevant representation. Subsequently, to deal with "overlapped multimodal representations", we propose Disentangling Learning Representations (DiLR) module to generate independent modality-unique and modality-common representation. The overall framework is shown in Fig. 2. A self-distillation strategy is used in training process, as illustrated in section 2.4.
@@ 3.2 Comparison with State-of-the-art Methods
We evaluate the performance of our model against three state-of-the-art multi-modality fusion techniques, as detailed in Table 1. For the purpose of comparison, we established a baseline method by utilizing Vision Transformer [4] and UNETR [5] as the backbones for the Fundus and OCT modality. The feature maps from both modalities are then simply concatenated and directly used for classification. The other compared methods include: (1) B-IF, which adopts an early fusion strategy; (2) M2LC [25], combining channel and spatial attention for enhanced feature integration; (3) IMDR [12], a strategy based on simple mutual information loss to implement cross-modality decoupling. The comparison is conducted in three settings, i.e., (1) complete modality, (2) one modality missing, (3) noisy modality.
Complete Modality Setting. In the ideal scenario withour any missing or noise, our model achieves the best performance among the models we test. Building upon this, we also test our approach under conditions where various Gaussian noise with different variance is introduced to each modality. As the noise level increases, a clear performance decline is observed in all models, emphasizing the challenges posed by data loss within a single modality on the stability of multimodal representations. Despite this, our method demonstrates exceptional robustness, particularly in scenarios with high levels of noise, consistently outperforming the other models.
@@ 3.3 Ablation Study
Effectiveness of each component. To evaluate the effectiveness of our novel EPRL and DiLR, we conducted an ablation study on the test set of Harvard-30k under gaussian noise with a variance of 0.5, as shown in Table 2. From Variant I to Variant II, we observe that the introduction of the EPRL unit effectively reduces the task-unrelated information within the modality, resulting in a significant increase in accuracy (6.5%). From Variant I to Variant III, we find that incorporating the DiLR unit effectively decouples modality-specific and modality-unique features, leading to an accuracy increase of 5%. In the final set of variants (Variant IV), where we combined both EPRL and DiLR, we observe that this configuration consistently outperforms both Variant II and Variant III, demonstrating that the two modules can work in synergy to provide highly decoupled and low-redundancy representations.
@@
Hyperparameter Sensitivity Analysis. To validate the robustness of our model, we conduct a series of hyperparameter sensitivity analysis in Table 3. In DiLR module, the percentage of common dimensions relative to total dimensions in modality embeddings may affect the performance. The results indicate that increasing shared information percentage during the early stages improves model performance. However, beyond a certain value, it may negatively impact the expression of specific information, leading to a decline in performance.

### v2

@@ Abstract
Ophthalmologists often rely on multimodal data to improve diagnostic precision. However, data on complete modalities are rare in real applications due to a lack of medical equipment and data privacy concerns. Traditional deep learning approaches usually solve these problems by learning representations in latent space. However, we highlight two critical limitations of these current approaches: (i) Task-irrelevant redundant information existing in complex modalities (e.g., massive slices) leads to a significant amount of redundancy in latent space representations. (ii) Overlapping multimodal representations make it challenging to extract features that are unique to each modality. To address these, we introduce the Essence-Point and Disentangle Representation Learning (EDRL) strategy that integrates a self-distillation mechanism into an end-to-end framework to enhance feature selection and disentanglement for robust multimodal learning. Specifically, Essence-Point Representation Learning module selects discriminative features that enhance disease grading performance. Moreover, the Disentangled Representation Learning module separates multimodal data into modality-common and modality-unique representations, reducing feature entanglement and enhancing both robustness and interpretability in ophthalmic disease diagnosis. Experiments on ophthalmology multimodal datasets demonstrate that the proposed EDRL strategy outperforms the state-of-the-art methods significantly. Code is available at GitHub Repository.
@@
footnotetext: *These authors contribute equallyfootnotetext: Corresponding authors: Zongyuan.Ge@monash.edu, Imran.razzak@mbzuai.ac.ae, Feilong.Tang@monash.edu
@@ 1 Introduction
To this end, we propose Essence-point and Disentangle Representation Learning (EDRL) framework. Specifically, the Essence-point Representation Learning (EPRL) module models the essence-points that guide the selection of discriminative information within each modality, effectively reducing task-irrelevant redundant representation. For modality feature disentanglement, the Disentangle Representation Learning (DiLR) module decomposes feature embeddings into modality-common and modality-unique components. It aligns shared information by driving the cross-correlation matrix toward a unit form while ensuring that modality-specific features remain distinct by approaching a zero form. Furthermore, we incorporate self-distillation between two pipelines (missing vs. complete modalities), using the complete pipeline to guide missing information reconstruction and enhance robustness under incomplete-modality conditions. Consequently, the EDRL framework minimizes redundancy, reduces inter-modality overlap, and enhances discriminative and generalizable multimodal learning.
Overall, our contributions are threefold. (i) We propose EPRL framework to reduce intra-modal redundancy by modeling prototypes for discriminative instance selection and incorporating self-distillation mechanism. (ii) We introduce DiLR to disentangle multimodal features into independent modality-unique and modality-common representation. (iii) Experiments on three ophthalmology multimodal datasets demonstrate the effectiveness of the EDRL strategy.
@@ 2.1 Problem Formulation
In this work, we propose an EDRL framework to address the challenge of missing modalities, particularly for solving (1) Inter-modality missing, where one modality is completely missing, and (2) Intra-modality missing, where natural noises are added into the modalities. To reduce "Task-irrelevant Redundant Information", we propose Essence-Point Representation Learning (EPRL) module to select the task-relevant representation. Subsequently, to deal with "overlapped multimodal representations", we propose Disentangling Learning Representations (DiLR) module to generate independent modality-unique and modality-common representation. The overall framework is shown in Fig. 2.
@@ 3.2 Comparison with State-of-the-art Methods
We evaluate the performance of our model against three state-of-the-art multi-modality fusion techniques, as detailed in Table 1. For the purpose of comparison, we established a baseline method by utilizing Vision Transformer [4] and UNETR [5] as the backbones for the Fundus and OCT modality. The feature maps from both modalities are then simply concatenated and directly used for classification. The other compared methods include: (1) B-IF, which adopts an early fusion strategy; (2) M2LC [25], combining channel and spatial attention for enhanced feature integration; (3) IMDR [12], a strategy based on simple mutual information loss to implement cross-modality decoupling. The comparison is conducted in three settings, i.e., (1) complete modality,(2) noisy modality (3) one modality missing.
Complete Modality and noisy modality Setting. In the ideal scenario without any missing or noise, our model achieves the best performance among the models we test. Building upon this, we also test our approach under conditions where various Gaussian noise with different variance is introduced to each modality (In Fig 3). As the noise level increases, a clear performance decline is observed in all models, emphasizing the challenges posed by data loss within a single modality on the stability of multimodal representations. Despite this, our method demonstrates exceptional robustness, particularly in scenarios with high levels of noise, consistently outperforming the other models.
@@ 3.3 Ablation Study
Effectiveness of each component. To evaluate the effectiveness of our novel EPRL and DiLR, we conducted an ablation study on the test set of Harvard-30k under gaussian noise with a variance of 0.5, as shown in Table 2. From Variant I to Variant II, we observe that the introduction of the EPRL unit effectively reduces the task-unrelated information within the modality, resulting in a significant increase in accuracy (6.5%). From Variant I to Variant III, we find that incorporating the DiLR unit effectively decouples modality-specific and modality-unique features, leading to an accuracy increase of 5%. Variant IV, combining both EPRL and DiLR, consistently outperforms Variants II and III, demonstrating their synergistic effect in producing highly decoupled, low-redundancy representations.
@@
Hyperparameter Sensitivity Analysis. To validate the robustness of our model, we conduct a series of hyperparameter sensitivity analysis in Table 3. In DiLR, the common dimension percentage affects performance: increasing it initially improves results, but excessive sharing impairs modality-specific information expression, causing performance decline.